Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

Actions

  • [Chandler]: Outline OSRO observation checklist (all required steps in delivery of archive data to PI). Changed to Gustaaf. This includes obtaining the output data characteristics (e.g., averaging) that are requested by the user.
  • [McMullin]: Add ECSO Projects to JIRA/tracking.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Describe heuristics for set-and-remember (describe latencies and proposed timing for set). Mark as closed; requirements are the same as for standard (non-band/frequency changing observations also require the dummy scan). This will continue to evolve but not as a separate category.
  • Proposed: [Perley]: Description of standard/required reduction steps in continuum processing. Accepted by Rick. Will track.

Commissioning review/reports

News

    • Michael, Martin, Dave away through Wednesday; holding current version steady (opportunity to review stability of the system). Very close to achieving the end-to-end automation but on-hold until they return. Joe, Vivek and Joan are scheduling and running the system (testimony to the hard work of the WIDAR team that this is possible; Ken supporting any issues)
    • Rick/Peggy returned from travel. Kumar back from travel.
    • Network switch went down over the weekend; being investigated
    • Observing suspended Monday morning due to snow in the dishes.
    • Executor updates planned for today. Also TelCal updates; should be invisible.
    • Half day shutdown due to HD upgrade (tomorrow?). Will happen Wednesday.

Notes for week of 15 Mar 2010 (Black: old but still relevant; Blue: new)

    • Due to issues of conflict/slow down for multiple access to the same file, please wait for John to copy to the data holding area at: /home/evla-data and grab it from there. This is happening now.
    • Project Codes should be getting into the system by early next week.
    • Production OPT has OSRO (but not RSRO) modes.
    • Many data sets taken over the weekend; under review (below)

OSRO Priority Testing Status

System Verification

      • High Frequency Readiness
        • Reference pointing (C/K: ECDATA-36): Report (Manuel). Appears to be working (other than for EA22). Would still like to have the test of a known offset and demonstrate that is recovered.
        • set-and-remember
          • within band frequency changes (ECDATA-33): okay (with antenna/subband exceptions).
            • Will see change until LO fiber is included in antenna phase model (Rick?)
        • Action: Rick/Vivek to develop final test this week to demonstrate readiness.
      • Gain transfer test (ECDATA-43): (Barry's test: src1-src2-src1-src1_called_1b-src1 - ...): Report (Vivek/Steve)*Evaluated and is working well (consistent dynamic range achieved for src and renamed src calibration transfers).*
      • L band, HI checkout (ECDATA-47): Report (Juergen)*Under review.*
      • Doppler setting Tests -Orion SiO at rise/set (ECDATA-48, ECDATA-41): Report (Lorant). Good news: Get same result at different times (this with Q band observing with no pointing used). Bad news: Velocity definition/reference frame are not making it into the SDM. Will need a new test when the scheduler is ready.
        • Action: Bryan will update on status of this.
      • Note effort to summarize tuning capabilities in: https://bugs.aoc.nrao.edu/browse/EVL-1137

Band Checkouts

      • Rick noted that he would like to add the frequency accuracy test (using known sources (beacons, RFI, etc). He will develop observations for this at each band.
        • L
          • Frequency labeling (ECDATA-9, ECDATA-10): Ok
          • Sky orientation (ECDATA-10): Ok
          • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
          • Tuning: being reviewed
          • RFI:
            • 1244, 1372: 1246, 1270, 1310, 1330, 1336-1338
            • 1350, 1810: Significant RFI in 1350
            • Action: Rick will provide an update to the default frequencies.
            • Note: Rick noted that there is a GPS signal (GPS L3) at 1385 that is there part of each day; this should be noted and watched for.
          • Antennas with no L-band: EA06, EA20, EA26
          • Integrating down (ECDATA-23):
            • Averaging in time is ok; averaging between subbands has problems.
            • L band (really) blank field (ECDATA-40):

        • S - Emmanuel
          • Frequency labeling (ECDATA-24): Ok
          • Sky orientation (ECDATA-18): Report (Emmanuel). Looks okay based on background sources in 'blank' field as compared to NVSS.
          • Gain stability (ECDATA-24): Ok (no jumps seen)
          • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
            • RFI: 3072 MHz, 3200 MHz (ECDATA-24)
            • Action: Rick will update the center frequency; will shift by 128 MHz for now; engineering fix is in the works.
            • Action: Rick will perform a frequency sweep.
          • Integrating down (ECDATA-18): Report (Emmanuel). On blank sky it integrates down.
          • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28
          • Note: Emmanuel noted that the known issue with having multiple spectral windows automatically filled causes problems in the data (currently this is done automatically so the data is corrupted). Bryan/Steve indicated that currently this just needs to be communicated as a warning to users that they must manually split out the data from the individual spectral windows.
            • Action: John will note when the archive will automatically do this correctly.
        • C
          • Frequency labeling: Ok. Will get more information from Eric's analysis.
          • Sky orientation (ECDATA-35): Ok
          • Gain stability (ECDATA-35): phase stable (30 min); amp jumps middle/between scans on EA06 D
          • Tuning: Low frequency tuning issue; stepped up for last AH1004.
            • RFI: Note: Rick noted that there is a microwave link across mountain tops which interferes with some antennas at 6020 MHz (with 40 MHz width). This must be tracked. It was noted that the OPT/Scheduler will try to guide observers away from known RFI sources.
          • Integrating down (ECDATA-25):
          • Antennas with no C-band: EA13
        • X (Nirupam, Ed)
          • Frequency labeling : ECDATA-17): Pending - Nirupam. Looks okay.
          • Sky orientation (ECDATA-17): Pending - Nirupam. Problems with the WICloop Offset data - Vivek will assist to see if an answer can be dug out. (post meeting update): Nirupam confirmed that the sky orientation was correct.
          • Gain stability (ECDATA-26): ok (no jumps)
          • Tuning: No issues
          • Integrating down (ECDATA-26): ok
          • Antennas
          • RFI: none seen
          • Note: Claire asked about the overall readiness for X band observing. There was a question about whether high end C band could be used in place of X band. This can be explored. Overall, if the sky orientation test is confirmed, we would like to push ahead with X band observing.
        • K (Claussen)
          • Frequency labeling ([https://bugs.aoc.nrao.edu/browse/ECDATA-19][ECDATA-19]]): ok
          • Sky orientation ([https://bugs.aoc.nrao.edu/browse/ECDATA-19][ECDATA-19]]): ok
          • Gain stability ([https://bugs.aoc.nrao.edu/browse/ECDATA-19][ECDATA-19]]): ok
          • Integrating down ([https://bugs.aoc.nrao.edu/browse/ECDATA-19][ECDATA-19]]): ok (close; differences seen when doing by channels).
          • RFI: none seen

Mode testing

      • Polarization (Myers/Moellenbrock); C band polcal run (ECDATA-12); follow-up observation (ECDATA-51) under review. L band polarization ECDATA-42: issues with SPW 0 due to interference. Report Myers/Moellenbrock
        • developed SB for a 5 hr run on 4 sources; 3C286 over transit (look at phase offset before and after transit. So far, C band looks okay. L band under review - complicated by lower subband having substantial RFI (causes Q, U to go berserk). Recommend to run through AIPS as well to ensure that users can process in this way (Action: Rick agreed to pass the polarization data through AIPS). Noted that OQ208 is not in AH10004 and doesn't have good parallactic angle coverage - should be inserted in if re-run.
      • Spectral Line - Bandpass stability: ECSV-35

OSRO Proposal Testing

    • AH1004; C band low data set taken; if okay, consider for release to PI. Report (Gustaaf)
    • AE176: review needed before more data taken; Report (Gustaaf/Jacqueline)*Result for scheduling tomorrow.*
    • AS983 (Trigger): Report (Vivek/Laura).*Complete*
    • AU129 (Blocks 2 and 4): Need review - needs assignment.*Gustaaf will assign.*
    • AM1009: Report (Emmanuel).*Under review (post-meeting update: the data are bad showing 16 sharp peaks across the BW; work to explore why this is happening - Ken asserts that these narrow channels had worked before).*
    • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • High Priority [Pokorny]: Auto-controlled CBE writes scrambled visibilities.
      • High Priority [Rupen]: X+ band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1).
      • Med Priority [TBD]: Delays in acquisition of source for some antennas. One antenna off source; EA23 is slow; not an issue.
      • Med Priority [Myers]: C band 128 MHz spectrum bends down at the edges.
      • Med Priority [Shores]: EA19-D station board won't boot.
      • Med Priority [Nick/Eric]: Update list of accepted names for calibrators.
      • Med Priority [Perley]: X band shows extra noise in set-and-remember test
      • Med Priority [TBD]: Subband 0 DC offset
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). EA03 R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Also seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history; new one seen (Dhawan, Myers) is 11s period; does not seem to degrade images too badly.

    • Lorant noted that he saw the same problem noted by Eric, Claire; 3 seconds of data from previous scan labeled as the other source). Add this to the list. Bryan asked for more information (essentially to note when this is happening and not happening to help understand the phenomenology and hunt down the solution. Bryan noted that the BB's time and the MCAF time are derived differently and so there is the possibility of an offset.

Commissioning plans for the week

System Tests

    • [Rick]: Enable High Frequency Observing: principally the reference pointing confirmation for C/K; understanding of the set-and-remember good/bad antennas to employ for calibration. If weather permits, test K band this week.
    • [Rick]: Enable RSRO Mode Observing
    • [Rupen; OSRO Priority 0]: OPT-> Archive path complete (currently just automated setup of the correlator).End of week possible
    • OSRO-1 (256 MHz): Still need tests of polarization phase convention
    • OSRO-2: Review fringes, gain/bandpass stability, phase sign, frequency labeling at L band.
    • OSRO-1,2: Review behavior under bandwidth options; change BW in OSRO 1 in steps of factor of 2 from 256 MHz -> 1 MHz and in OSRO 2 in steps of factor of 2 from 256 MHz -> 250 kHz

OSRO project tests

As possible; see overview at: https://staff.nrao.edu/wiki/bin/view/EVLA/OSROReadiness_Cycle1; Note: Once automated corr setup is enabled; open up the ECSV plan observations (see planning review).
  • Tool from Lorant to see available SBs: /home/mchost/evla/bin/sbStatReport
    • AE176
    • AU129
    • AM1009
    • AR704
    • AV317

RSRO

  • Observations
  • Residents
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Adam Deller (POC: Bryan Butler; area(s): Correlator checkout)
    • James Miller-Jones (POC: Claire Chandler; area(s): Demo Science)
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests)
    • Dave Green (arriving; POC: Sanjay Bhatnagar); area(s): Imaging

This week in JIRA

  • JIRA and e-mail notifications

Lots of ways to generate e-mail
  • Currently the reporter/assignee receive all e-mail traffic
  • In addition, a core group of system verification/OSRO testers (ECDATA Group) also receive all of the following notifications; this group includes:
    • Joe, Rick, Michael, Steve, Vivek, Gustaaf, Bryan, Emmanuel, Juergen, Mark

Current and Modified JIRA notifications include:
Category Event Reporter/Assignee ECDATA Group Watchers Notes
Issue        
  Created x x x needed
  Updated x x x needed
  Assigned x x x removed for ECDATA Group/Watchers
  Resolved x x x needed
  Closed x x x needed
  Issue Reopened x x x needed
  Issue Deleted x x x removed for ECDATA Group/Watchers
  Issue Moved x x x removed for ECDATA Group
Comments        
  Commented x x x needed
  Comment Edited x x x needed
Worklogs       Worklogs not used for EVLA
  Work logged on Issue x x x removed for ECDATA Group/Watchers
  Work started on issue x x x removed for ECDATA Group/Watchers
  Work stopped on issue x x x removed for ECDATA Group/Watchers
  Issue worklog updated x x x removed for ECDATA Group/Watchers
  Issue worklog deleted x x x removed for ECDATA Group/Watchers
Generic Event - x x x needed

  • What's the difference between resolved and closed?
    • The intention is that the person assigned to the action/task is able to resolve it (based on the description of what is required). It is intended to be closed by the submitter to indicate that the resolving solution satisfies the full intent of the submitted action.

The data acquisition rate has gone up substantially requiring more people to be involved in the reviews; working well so far but not clear how well we'll scale. Ultimately, each OSRO/RSRO/ECSO projects won't need to be specifically tested and so this is a temporary situation based on the state of commissioning.

  • Given the amount of data, how do I find what I'm looking for?

AOB

  • Chris asked whether C band for reference pointing was the recommendation; yes for now. Frazer asked whether using C instead of X changed the timing of the observations substantially; Rick, no, they're next to each other so 2.5 minutes for pointing should be fine (but we should check this).
  • Rick asked why the TP is sampled only 1/minute. Bryan noted that is what the engineers wanted but it can be modified/monitored at a different rate as needed.

Planning review (for Q1/Q2 2010):

Followup deferred. Begin to look at the detailed needs for the tests being scoped out by the leads. Want to finalize the detailed observing/testing plans by the end of this month and have a view of the longer term commitments needed. Establish an overview of the obstacles to deploying to the community; in particular, understand all of the details of the end-to-end data flow/processing to make decisions on what is working versus what is usable by the community. Something like:

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
Example limit of 1 s int ok OPT selected limits not possible no recommendations avail ok can't cook coffee weather table - transient issue no-go

  • Critical Commissioning
    • System Verification (Perley/Rupen): Need plan to conclude complex gain, bandpass stability tests at L-Q, reference pointing, gain curves, tipping scans.
    • Observing Band Checkout (principally calibration strategies, flagging)
    • Observing Mode Support (broad modes based on science goals)
    • Observing support (e2e capability)
      • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
      • Flux Calibrator Models (Amy); ECSV-43
  • Advanced Science Support

    • Need:
      • [Butler] Planning: Planetary observing (OPT setup; correlator needs); ECSV-7, ECSV-15
      • [Chandler] Planning: Demo Science
      • [Frail] Planning: ToO scheduling heuristics
      • [Sjouwerman] Planning: OPT tests for primary frequencies; ECSV-6
      • [van Moorsel] Planning: Data distribution plan
      • [Sjouwerman] Planning: Archive functionality tests; ECSV-11
      • [Mioduszewski] Planning: Calibrator survey; ECSV-27
      • [Brisken] Planning: Phased array (VLBI, pulsars); ECSV-22, ECSV-23
      • [Golap] Planning CASA multi-thread/cluster testing.
      • [Walker] Planning VLBI setup

-- JosephMcMullin - 15 Mar 2010
Topic revision: r4 - 2010-03-16, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback